Operations Report – November 15th

Crew 182 Operations Report  15Nov2017

SOL: 12

Name of person filing report:  Carmen Atauconcha

Non-nominal systems:

WATER HEATER: Finally, we found the problem. The new water heater is not installed yet.

CHARGERS OF THE SUITS: Two of the chargers are working, but we moved them and now it is hard for us to connect them, again.


Generator—  It worked the whole night since 17:05pm of 14Nov2017 until 8:00 am of 15Nov2017.  Turn on at 16:30 pm 15Nov2017.  Operand.




9:30 am 100%
10:30 am 100%
11:30 am 100%
12:30 pm 100%
13:30 pm 100%
14:35 pm 98%
15:30 pm 90/%
16:00 pm 85%
16:30 pm 80%
17:30 pm 100%


Diesel – 50 %

Propane – 49 psi

Gasoline (5 Gallon containers for ATV) –  2 ½ gallons.

Water (trailer) –  0 gallons

Water (static) – 350  gallons

Trailer to Static Pump used –   No

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  62 gallons

Toilet tank emptied:  yes

ATV’s Used: Yes

Oil Checked:  No

Atv # Fuel Used Gals:  1  gallons

Tires Status: Perfect

# Hours the ATVs were Used today: 2 ½  hours

Deimos used:  Si

Hours: 2 ½

Beginning charge:  – 14:20

Ending charge:  –

Currently charging: Yes
HabCar used? No

Notes and Comments:

We love the new ATV’s and Rovers. Thanks¡

Summary of Hab Operations:
In the morning, when we waked up, we received four more  ATV’s. It is nice, of all them work. Also today, we went to recover the ATV left yesterday in the desert. Also, we needed to charge it, because it does not have battery. We connected it with another ATV, and it works.

Summary of GreenHab operations:

We are cleaning and organizing this area. It is a disaster, but we are working on it.

Summary of ScienceDome operations:


We are using the lamps of the Hab, for the laboratory.

Summary of health and safety issues:

Atila is better now.  All of us are better that yesterday.

Questions, concerns and requests to Mission Support:

We need gasoline for the ATV’S, please. URGENT

Copyright © The Mars Society. All rights reserved. | Main Site